A chemist has one solution that is 25% salt and 75% water, and another solution that is 10% salt. How many liters of each sould the chemist use to make 120 liters of a solution that is 15% salt?
Answer by stanbon(75887)

A chemist has one solution that is 25% salt and 75% water, and another solution that is 10% salt. How many liters of each sould the chemist use to make 120 liters of a solution that is 15% salt?
-----------------------
25% solution data:
Amt. = x liters ; amt of salt = x/4 liters
--------------------
10% solution data:
Amt = 120-x liters ; amt of salt = 0.1(120-x) = 12-0.1x liters
--------------------
15% solution data:
Amt = 120 liters ; amt of salt = 0.15*120 = 18 liters
----------------------
EQUATION:
salt + salt = salt
x/4 + 12-0.1x = 18
Multiply thru by 40 to get:
10x + 480 - 4x = 18*40
14x = 18*40-480
x = 17.14 liters (amount of 25% solution in the mixture)
120-x = 102.86 liters (amount of 10% solution in the mixture)
================
Cheers,
Stan H.